# Non-uniqueness of the steady state for run-and-tumble particles with a double-well interaction potential

### v4

@@ Abstract
We study $N$ run-and-tumble particles (RTPs) in one dimension interacting via a double-well pairwise potential $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which is repulsive at short interparticle distance $r$ and attractive at large distance. At large time, the system forms a bound state where the density of particles has a finite support. We focus on the determination of the total density of particles in the stationary state $\rho_{s}(x)$, in the limit $N\to+\infty$. We obtain an explicit expression for $\rho_{s}(x)$ as a function of the “renormalized” interaction parameter $k=k_{0}-3m_{2}$ where $m_{2}$ is the second moment of $\rho_{s}(x)$. Interestingly, this stationary solution exhibits a transition between a connected and a disconnected support for a certain value of $k$, which has no equivalent in the case of Brownian particles. Analyzing in detail the expression of the stationary density in the two cases, we find a variety of regimes characterized by different behaviors near the edges of the support and around $x=0$. Furthermore, by studying the relation between $k$ and $k_{0}$, we find that the mapping $k_{0}\to k$ becomes multi-valued below a certain value of the tumbling rate $\gamma$ of the RTPs for some range of values of $k_{0}$ near the transition, implying the existence of two stable solutions. Finally, we show that in the case of a disconnected support, it is possible to observe steady states where the density $\rho_{s}(x)$ is not symmetric, characterized by a third moment $m_{3}$ which can take a continuous range of values. All our analytical predictions are in good agreement with numerical simulations already for systems of $N=100$ particles. The non-uniqueness of the stationary state is a particular feature of this model in the presence of active (RTP) noise, which contrasts with the uniqueness of the Gibbs equilibrium for Brownian particles. We argue that these results are also relevant for a class of more realistic interactions with both an attractive and a repulsive part, but which decay at infinity.
@@ I.1 Overview
In the continuum, one of the simplest models of active particle is the run-and-tumble particle (RTP), inspired by the motion of E. Coli bacteria Berg2004 . In one dimension, it is driven by so-called telegraphic noise, which alternates between two values at a constant rate HJ95 ; W02 ; ML17 ; kac74 ; Orshinger90 ; Schnitzer ; TailleurCates . Recently we have considered a system of $N$ RTPs in one dimension, interacting via a pairwise power law potential RieszFluct . In particular we have studied two cases of long range interactions, the logarithmic potential, which can be seen as an active generalization of the Dyson Brownian motion TouzoDBM2023 , and the linear 1D Coulomb potential, also called active rank diffusions activeRDshort ; activeRDlong . In both cases, the aim was to compute the density of particles in the steady state in the limit of large $N$. In the case of an attractive 1D Coulomb interaction, one finds that the particles form a bound state which exhibits a transition between a smooth density with unbounded support, and a density with bounded support displaying clusters at the edges activeRDshort ; activeRDlong .
@@
In this paper we consider a toy model for such interactions, which consists in $N$ one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique. We note that, although this work focuses on the large $N$ limit of the model, the existence of multiple steady states and the breaking of parity symmetry are already visible for small values of $N$, as we have observed from numerical simulations, which we report in Appendix B.
Our model has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large $N$ limit, derived in activeRDshort , can be solved analytically. Furthermore, although our model has an interaction force $W^{\prime}(r)$ which diverges at large distance, which is a priori an unrealistic feature, one can make an argument that a similar behavior can be observed for more realistic interaction forces which vanish at infinity, but for which the system still has a bound state. A more precise criterion for the existence of a bound state and for the transition from single to disjoint support is discussed in Appendix C, and is confirmed by a numerical simulation. In that case we expect all the results obtained for the double well toy model to be qualitatively relevant.
@@ I.2 Model and methodology
We consider $N$ run-and-tumble particles in one dimension interacting via a two-body potential $W(x)$, which can be described by the equations of motion
@@
If $W^{\prime}(0)=0$, the particles can always cross even for $T=0$, and one can use the Dean-Kawasaki method Dean ; Kawa for 1D RTPs introduced in TouzoDBM2023 . In the large $N$ limit, for $T=0$, we have shown in activeRDshort (the derivation is recalled in Appendix A) that the stationary density obeys the self-consistent equation
@@ I.3 Main results
    where $A$ is a normalization constant. From this one can deduce the critical value of $k_{0}$ which corresponds to $k=k_{c}=3/2^{2/3}$,
@@
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed direct numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady states (see Fig. 8) are confirmed by these numerical results. The case of small values of $N$ is discussed in Appendix B.
@@ II.1 Passive case
for any (sufficiently attractive) interaction potential $W$ such that $\cal P$} is normalizable on the real axis. In the limit of large $N$ one can describe the system by a density field $\rho(x,t)=\frac{1}{N}\sum_{i}\delta(x-x_{i})$, with a probability distribution (see e.g. DeanMajumdar2008 )
@@
where the entropy term (second term) arises from the Jacobian of the mapping between the two descriptions, and $i\lambda$ is a Lagrange multiplier enforcing the constraint $\int dx\rho(x)=1$. In the large $N$ limit the functional integral is dominated by a saddle point. Taking a functional derivative with respect to the density of the term in the exponential, we obtain that the optimal density satisfies
@@ II.2 Calculation of the stationary density in the purely active case
Finally, we have tested these predictions using numerical simulations. We solve numerically the equation of motion (1) at $T=0$ for $N=100$ particles, starting from a uniform initial condition on some interval $x_{init},x_{init}]$ and we determine the stationary total density by time averaging the density in the center of mass frame at large time. For $k>k_{c}$, we instead start from an initial condition where the particles are separated into two groups of equal size, which allows to obtain a symmetric stationary state. Different choices of initial condition may lead to an asymmetric stationary state, see Sec. III. The results are displayed in Fig. 3 and show excellent agreement (although finite $N$ effects are visible near the edges of the support in some cases).
@@ II.3.3 Bistability
For $N\to+\infty$ and $\gamma<\gamma_{c}$, there exists an interval of values of $k_{0}$ around $k_{0,c}$ where there are three possible values of kk [^1] Note that for a given value of $k$ there is however a unique corresponding value of $k_{0}$, hence in particular the critical value $k_{0,c}=k_{0}(k_{c})$ is uniquely defined., see Fig. 4 (bottom). There are two possible situations. (i) For most values of $\gamma$ and $k_{0}$, two of these stationary state have a connected support ($k<k_{c}$) and the other one has a disconnected support ($k>k_{c}$). (ii) In a small region of parameters $k_{0}$ and $\gamma$ however, we have noticed that all three solutions correspond to $k<k_{c}$, i.e. to a connected support, see Fig. 4 (bottom right). All three values correspond a priori to possible stationary states of the system, which however can be stable or unstable. We have not studied analytically the stability of these solutions (which seems a priori technically challenging). However, we have compared these predictions with numerical simulations for finite $N$ as we now discuss.
@@
Note that an important consequence of this bistability is that in some range of parameters the transition between a connected and disconnected support is now discontinuous as a function of $k_{0}$. Finally, there are two interesting questions for future investigations. It would be nice to be able to prove the stability properties of the different stationary solutions at infinite $N$. It is an open question whether at finite $N$ the system can transition from one stationary state to the other over an extremely long period of time.
@@ III Non symmetric steady state
Until now we have assumed that the stationary density $\rho_{s}(x)$ is symmetric, $\rho_{s}(-x)=\rho_{s}(x)$. This should be true in the large $N$ limit if we start from a symmetric initial condition. It should also hold whenever the support has a single component, due to ergodicity. However, as we now show, if the initial condition is not symmetric, and if the support is disjoint, a non-symmetric stationary distribution is possible, where the particles are separated into two groups of different sizes. Such a distribution is characterized by a non-zero value of the third moment $m_{3}$.
@@ III.2 Detailed solution and support of the density
As in the symmetric case, one can determine a posteriori the corresponding value of $k_{0}=k+3m_{2}$. Note that $m_{2}$ now depends on $m_{3}$, so that the mapping between $k$ and $k_{0}$ is different for each value of $m_{3}$ (and in particular it differs from the symmetric case $m_{3}=0$). By contrast, the second “effective” parameter $m_{3}$, which controls the asymmetry of the solution, cannot be mapped directly to one of the “true” parameters of the model. One can compute a posteriori the proportion $m_{0}^{R}$ of particles on the right side of the support, which is another, perhaps more intuitive way to quantify the asymmetry of the steady state. However, this asymmetry is determined in a non-trivial way by the initial condition, as well as by the history of the noise in the case of large but finite $N$ that we consider in the simulations (see below). The detailed study of this dependence goes beyond the scope of this paper.
@@ III.4 Numerical results
We have tested some of the predictions above in numerical simulations (for $N=100$ particles). One finds that indeed starting from an asymmetric initial condition, e.g. separating the particles into two groups of unequal size at time $t=0$, one can obtain a non-symmetric stationary density, which is well described by the expression (78)-(79) – see Fig. 8 for an example. In addition we have studied numerically small values of $N$ in Appendix B. Interestingly, non-symmetric stationary densities can already be observed for $N=5$. However, they did not arise in our simulations for $N=3$ or $4$, which could be interpreted as a sign that the limiting value $\alpha=1/3$ computed above for $N\to+\infty$ also holds at finite $N$.
@@ IV Conclusion
In this paper we have considered $N$ run-and-tumble particles in one dimension interacting via a pairwise double well potential $W(r)$, with associated parameter $k_{0}=-W^{\prime\prime}(0)$. Thanks to a self consistent equation valid in the large $N$ limit, we were able to express the stationary density $\rho_{s}(x)$ in the large $N$ limit as a function of a renormalized parameter $k=k(k_{0})$, and of the tumbling rate $\gamma$. We showed that $\rho_{s}(x)$ exhibits a transition, where the particles separate into two groups for $k>k_{c}$. In both phases the density displays power law singularities at the edges of the support. When the two groups are of the same size, we find that in some range of parameters around the transition, $k(k_{0})$ is multivalued, which implies the existence of two stable steady states. Furthermore, we show that for $k>k_{c}$ it is possible to observe non-symmetric (non parity invariant) steady states which are characterized by a fraction $\alpha\neq 1/2$ of particles in the rightmost group. We show that $\alpha$ can vary continuously inside an interval that we determine. All these analytical predictions, valid for $N\to+\infty$, have been compared with numerical simulations for $N=100$, with very good agreement. In particular we have checked that different steady states can be reached by changing the initial condition.
@@
There are several open problems even for this simple type of model. For instance one would like to be able to solve for the dynamics and obtain the basin of attraction of each steady state. Another question is whether, for finite $N$, the noise can allow for rare transitions between these steady states, and at which rate. One could also try to extend this to higher order degree polynomials (or more general functions) and determine the co-dimension of the stationary manifold.
@@ Appendix A Self-consistent equation for the stationary density
Let us recall briefly the derivation of the self-consistent equation (4) for the stationary density in the large $N$ limit in the case of the RTPs activeRDshort , in the absence of passive noise, i.e. for $T=0$ but with an additional external potential $V(x)$. We first consider a general interaction potential $W(x)$, with $W^{\prime}(0)=0$, such that the particles can cross. In this case, one can use the Dean-Kawasaki method Dean ; Kawa extended to the RTPs TouzoDBM2023 to obtain the evolution equation for the densities $\rho_{\sigma}$ with $\sigma=\pm 1$ in the limit of large $N$,
@@ Appendix B Double-well interaction for $N=2$, and numerics for small values of NN
The support of $X$ can be obtained by studying the fixed points of this equation of motion. This is very similar to the case $N\to+\infty$ (see Fig. 2), but with a non-renormalized $k_{0}$. In addition to the two lines $f(X)=\pm 2v_{0}$, there are additional fixed points corresponding to $f(X)=0$ (corresponding to the $\sigma_{1}=\sigma_{2}$), but they do not play any role in determining the support. Thus there is a transition for some values of $k_{0}$ between a regime where $X$ can take all values inside an interval $X_{e},X_{e}]$, meaning that the two particles can cross, and a state where they cannot cross, i.e. the support of $X$ is disjoint. This transition occurs when the local maximum at $X^{*}=\sqrt{\frac{k_{0}}{3g}}$ reaches the value $2v_{0}$, i.e. when $2\left(\frac{k_{0}}{3g}\right)^{3/2}=v_{0}$, or $k_{0}=3g(v_{0}/2)^{2/3}$ (replacing $k_{0}$ by $k$ and taking $g=1$ and $v_{0}=1$ we recover $k_{c}$). For $N=2$ this transition is rather straightforward, but the fact that it persists at large $N$ is not a priori obvious. Note that for $N=2$ the particle density is always symmetric around the center of mass and thus $m_{3}$ is necessarily zero in this case. The symmetry breaking discussed in the text is thus a property which appears for larger values of $N$.
@@
These results show that the existence of multiple stationary states is not just an effect of large $N$ but is also present for small values of $N$ (e.g. for $N=5$ one may have either $2$ particles on the left and $3$ on the right or vice versa).

### v5

@@ Abstract
We study NN run-and-tumble particles (RTPs) in one dimension interacting via a double-well pairwise potential $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which is repulsive at short interparticle distance $r$ and attractive at large distance. At large time, the system forms a bound state where the density of particles has a finite support. We focus on the determination of the total density of particles in the stationary state $\rho_{s}(x)$, in the limit $N\to+\infty$. We obtain an explicit expression for $\rho_{s}(x)$ as a function of the “renormalized” interaction parameter $k=k_{0}-3m_{2}$ where $m_{2}$ is the second moment of $\rho_{s}(x)$. Interestingly, this stationary solution exhibits a transition between a connected and a disconnected support for a certain value of $k$, which has no equivalent in the case of Brownian particles. Analyzing in detail the expression of the stationary density in the two cases, we find a variety of regimes characterized by different behaviors near the edges of the support and around $x=0$. Furthermore, by studying the relation between $k$ and $k_{0}$, we find that the mapping $k_{0}\to k$ becomes multi-valued below a certain value of the tumbling rate $\gamma$ of the RTPs for some range of values of $k_{0}$ near the transition, implying the existence of two stable solutions. Finally, we show that in the case of a disconnected support, it is possible to observe steady states where the density $\rho_{s}(x)$ is not symmetric, characterized by a third moment $m_{3}$ which can take a continuous range of values. All our analytical predictions are in good agreement with numerical simulations already for systems of $N=100$ particles. The non-uniqueness of the stationary state is a particular feature of this model in the presence of active (RTP) noise, which contrasts with the uniqueness of the Gibbs equilibrium for Brownian particles. We argue that these results are also relevant for a class of more realistic interactions with both an attractive and a repulsive part, but which decay at infinity.
@@ I.1 Overview
In the continuum, one of the simplest models of active particle is the run-and-tumble particle (RTP), inspired by the motion of E. Coli bacteria Berg2004 . In one dimension, it is driven by so-called telegraphic noise, which alternates between two values at a constant rate HJ95 ; W02 ; ML17 ; kac74 ; Orshinger90 ; Schnitzer ; TailleurCates . Recently we have considered a system of NN RTPs in one dimension, interacting via a pairwise power law potential RieszFluct . In particular we have studied two cases of long range interactions, the logarithmic potential, which can be seen as an active generalization of the Dyson Brownian motion TouzoDBM2023 , and the linear 1D Coulomb potential, also called active rank diffusions activeRDshort ; activeRDlong . In both cases, the aim was to compute the density of particles in the steady state in the limit of large NN. In the case of an attractive 1D Coulomb interaction, one finds that the particles form a bound state which exhibits a transition between a smooth density with unbounded support, and a density with bounded support displaying clusters at the edges activeRDshort ; activeRDlong .
@@
In this paper we consider a toy model for such interactions, which consists in NN one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique. We note that, although this work focuses on the large NN limit of the model, the existence of multiple steady states and the breaking of parity symmetry are already visible for small values of NN, as we have observed from numerical simulations, which we report in Appendix B.
Our model has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large NN limit, derived in activeRDshort , can be solved analytically. Furthermore, although our model has an interaction force $W^{\prime}(r)$ which diverges at large distance, which is a priori an unrealistic feature, one can make an argument that a similar behavior can be observed for more realistic interaction forces which vanish at infinity, but for which the system still has a bound state. A more precise criterion for the existence of a bound state and for the transition from single to disjoint support is discussed in Appendix C, and is confirmed by a numerical simulation. In that case we expect all the results obtained for the double well toy model to be qualitatively relevant.
@@ I.2 Model and methodology
We consider NN run-and-tumble particles in one dimension interacting via a two-body potential $W(x)$, which can be described by the equations of motion
@@
If $W^{\prime}(0)=0$, the particles can always cross even for $T=0$, and one can use the Dean-Kawasaki method Dean ; Kawa for 1D RTPs introduced in TouzoDBM2023 . In the large NN limit, for $T=0$, we have shown in activeRDshort (the derivation is recalled in Appendix A) that the stationary density obeys the self-consistent equation
@@ I.3 Main results
    where AA is a normalization constant. From this one can deduce the critical value of $k_{0}$ which corresponds to $k=k_{c}=3/2^{2/3}$,
@@
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed direct numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady states (see Fig. 8) are confirmed by these numerical results. The case of small values of NN is discussed in Appendix B.
@@ II.1 Passive case
for any (sufficiently attractive) interaction potential $W$ such that $\cal P$} is normalizable on the real axis. In the limit of large NN one can describe the system by a density field $\rho(x,t)=\frac{1}{N}\sum_{i}\delta(x-x_{i})$, with a probability distribution (see e.g. DeanMajumdar2008 )
@@
where the entropy term (second term) arises from the Jacobian of the mapping between the two descriptions, and $i\lambda$ is a Lagrange multiplier enforcing the constraint $\int dx\rho(x)=1$. In the large NN limit the functional integral is dominated by a saddle point. Taking a functional derivative with respect to the density of the term in the exponential, we obtain that the optimal density satisfies
@@ II.2 Calculation of the stationary density in the purely active case
Finally, we have tested these predictions using numerical simulations. We solve numerically the equation of motion (1) at $T=0$ for $N=100$ particles, starting from a uniform initial condition on some interval $x_{init},x_{init}]$ and we determine the stationary total density by time averaging the density in the center of mass frame at large time. For $k>k_{c}$, we instead start from an initial condition where the particles are separated into two groups of equal size, which allows to obtain a symmetric stationary state. Different choices of initial condition may lead to an asymmetric stationary state, see Sec. III. The results are displayed in Fig. 3 and show excellent agreement (although finite NN effects are visible near the edges of the support in some cases).
@@ II.3.3 Bistability
For $N\to+\infty$ and $\gamma<\gamma_{c}$, there exists an interval of values of $k_{0}$ around $k_{0,c}$ where there are three possible values of kk [^1] Note that for a given value of $k$ there is however a unique corresponding value of $k_{0}$, hence in particular the critical value $k_{0,c}=k_{0}(k_{c})$ is uniquely defined., see Fig. 4 (bottom). There are two possible situations. (i) For most values of $\gamma$ and $k_{0}$, two of these stationary state have a connected support ($k<k_{c}$) and the other one has a disconnected support ($k>k_{c}$). (ii) In a small region of parameters $k_{0}$ and $\gamma$ however, we have noticed that all three solutions correspond to $k<k_{c}$, i.e. to a connected support, see Fig. 4 (bottom right). All three values correspond a priori to possible stationary states of the system, which however can be stable or unstable. We have not studied analytically the stability of these solutions (which seems a priori technically challenging). However, we have compared these predictions with numerical simulations for finite NN as we now discuss.
@@
Note that an important consequence of this bistability is that in some range of parameters the transition between a connected and disconnected support is now discontinuous as a function of $k_{0}$. Finally, there are two interesting questions for future investigations. It would be nice to be able to prove the stability properties of the different stationary solutions at infinite NN. It is an open question whether at finite NN the system can transition from one stationary state to the other over an extremely long period of time.
@@ III Non symmetric steady state
Until now we have assumed that the stationary density $\rho_{s}(x)$ is symmetric, $\rho_{s}(-x)=\rho_{s}(x)$. This should be true in the large NN limit if we start from a symmetric initial condition. It should also hold whenever the support has a single component, due to ergodicity. However, as we now show, if the initial condition is not symmetric, and if the support is disjoint, a non-symmetric stationary distribution is possible, where the particles are separated into two groups of different sizes. Such a distribution is characterized by a non-zero value of the third moment $m_{3}$.
@@ III.2 Detailed solution and support of the density
As in the symmetric case, one can determine a posteriori the corresponding value of $k_{0}=k+3m_{2}$. Note that $m_{2}$ now depends on $m_{3}$, so that the mapping between $k$ and $k_{0}$ is different for each value of $m_{3}$ (and in particular it differs from the symmetric case $m_{3}=0$). By contrast, the second “effective” parameter $m_{3}$, which controls the asymmetry of the solution, cannot be mapped directly to one of the “true” parameters of the model. One can compute a posteriori the proportion $m_{0}^{R}$ of particles on the right side of the support, which is another, perhaps more intuitive way to quantify the asymmetry of the steady state. However, this asymmetry is determined in a non-trivial way by the initial condition, as well as by the history of the noise in the case of large but finite NN that we consider in the simulations (see below). The detailed study of this dependence goes beyond the scope of this paper.
@@ III.4 Numerical results
We have tested some of the predictions above in numerical simulations (for $N=100$ particles). One finds that indeed starting from an asymmetric initial condition, e.g. separating the particles into two groups of unequal size at time $t=0$, one can obtain a non-symmetric stationary density, which is well described by the expression (78)-(79) – see Fig. 8 for an example. In addition we have studied numerically small values of NN in Appendix B. Interestingly, non-symmetric stationary densities can already be observed for $N=5$. However, they did not arise in our simulations for $N=3$ or 44, which could be interpreted as a sign that the limiting value $\alpha=1/3$ computed above for $N\to+\infty$ also holds at finite NN.
@@ IV Conclusion
In this paper we have considered NN run-and-tumble particles in one dimension interacting via a pairwise double well potential $W(r)$, with associated parameter $k_{0}=-W^{\prime\prime}(0)$. Thanks to a self consistent equation valid in the large NN limit, we were able to express the stationary density $\rho_{s}(x)$ in the large NN limit as a function of a renormalized parameter $k=k(k_{0})$, and of the tumbling rate $\gamma$. We showed that $\rho_{s}(x)$ exhibits a transition, where the particles separate into two groups for $k>k_{c}$. In both phases the density displays power law singularities at the edges of the support. When the two groups are of the same size, we find that in some range of parameters around the transition, $k(k_{0})$ is multivalued, which implies the existence of two stable steady states. Furthermore, we show that for $k>k_{c}$ it is possible to observe non-symmetric (non parity invariant) steady states which are characterized by a fraction $\alpha\neq 1/2$ of particles in the rightmost group. We show that $\alpha$ can vary continuously inside an interval that we determine. All these analytical predictions, valid for $N\to+\infty$, have been compared with numerical simulations for $N=100$, with very good agreement. In particular we have checked that different steady states can be reached by changing the initial condition.
@@
There are several open problems even for this simple type of model. For instance one would like to be able to solve for the dynamics and obtain the basin of attraction of each steady state. Another question is whether, for finite NN, the noise can allow for rare transitions between these steady states, and at which rate. One could also try to extend this to higher order degree polynomials (or more general functions) and determine the co-dimension of the stationary manifold.
@@ Appendix A Self-consistent equation for the stationary density
Let us recall briefly the derivation of the self-consistent equation (4) for the stationary density in the large NN limit in the case of the RTPs activeRDshort , in the absence of passive noise, i.e. for $T=0$ but with an additional external potential $V(x)$. We first consider a general interaction potential $W(x)$, with $W^{\prime}(0)=0$, such that the particles can cross. In this case, one can use the Dean-Kawasaki method Dean ; Kawa extended to the RTPs TouzoDBM2023 to obtain the evolution equation for the densities $\rho_{\sigma}$ with $\sigma=\pm 1$ in the limit of large NN,
@@ Appendix B Double-well interaction for $N=2$, and numerics for small values of NN
The support of $X$ can be obtained by studying the fixed points of this equation of motion. This is very similar to the case $N\to+\infty$ (see Fig. 2), but with a non-renormalized $k_{0}$. In addition to the two lines $f(X)=\pm 2v_{0}$, there are additional fixed points corresponding to $f(X)=0$ (corresponding to the $\sigma_{1}=\sigma_{2}$), but they do not play any role in determining the support. Thus there is a transition for some values of $k_{0}$ between a regime where $X$ can take all values inside an interval $X_{e},X_{e}]$, meaning that the two particles can cross, and a state where they cannot cross, i.e. the support of $X$ is disjoint. This transition occurs when the local maximum at $X^{*}=\sqrt{\frac{k_{0}}{3g}}$ reaches the value $2v_{0}$, i.e. when $2\left(\frac{k_{0}}{3g}\right)^{3/2}=v_{0}$, or $k_{0}=3g(v_{0}/2)^{2/3}$ (replacing $k_{0}$ by $k$ and taking $g=1$ and $v_{0}=1$ we recover $k_{c}$). For $N=2$ this transition is rather straightforward, but the fact that it persists at large NN is not a priori obvious. Note that for $N=2$ the particle density is always symmetric around the center of mass and thus $m_{3}$ is necessarily zero in this case. The symmetry breaking discussed in the text is thus a property which appears for larger values of NN.
@@
These results show that the existence of multiple stationary states is not just an effect of large NN but is also present for small values of NN (e.g. for $N=5$ one may have either 22 particles on the left and 33 on the right or vice versa).